OpenStudy (anonymous):

An object is dropped from a 400-ft tower. a) when does it hit the ground? b) how fast is it going at the time of impact?

OpenStudy (anonymous):

the acceleration due to gravity is 32 ft/s^2. \[a(t)=f^{''}(t)=-32\] \[v(t)=f^{'}(t)=-32t+c\] \[d(t)=f(t)=-16t^2+ct+k\] given we are dropping the object from a height of 400ft \[v(0)=0 \rightarrow c=0; d(0)=400 \rightarrow k=400\] \[v(t)=-32t; d(t)=400-16t^2\] when it hits the ground: \[0=400-16t^2; t=\pm5\] although -5 is technically true it does not make sense in the problem \[t=5 \rightarrow v(5)=-160\] so a)= 5 seconds b)= -160 ft/s